### v3

# Alexandra Ciobotaru

Abstract
Psychological corpora in NLP are collections of texts used to analyze human psychology, emotions, and mental health. These texts allow researchers to study psychological constructs, detect mental health issues and analyze emotional language. However, mental health data can be difficult to collect correctly from social media, due to suppositions made by the collectors. A more pragmatic strategy involves gathering data through open-ended questions and then assessing this information with self-report screening surveys. This method was employed successfully for English, a language with a lot of psychological NLP resources. However, this cannot be stated for Romanian, which currently has no open-source mental health corpus. To address this gap, we have created the first open-source corpus for depression and anxiety in Romanian, by utilizing a form with 6 open-ended questions along with the standardized PHQ-9 and GAD-7 screening questionnaires. Consisting of the texts of 205 respondents and although it may seem small, PsihoRo is a first step towards understanding and analyzing texts regarding the mental health of the Romanian population. We employ statistical analysis, text analysis using Romanian LIWC, emotion detection and topic modeling to show what are the most important features of this newly introduced resource to the NLP community. The data is publicly available at https://huggingface.co/datasets/Alegzandra/PsihoRo.
Keywords mental health $\cdot$ Romanian $\cdot$ emotion detection $\cdot$ LIWC $\cdot$ topic modeling
1 Introduction
Mental health research has increasingly used data from social media platforms, which often present a curated view of individuals’ personal lives. While social media can provide valuable insights for mental health studies [24, 30], there is a risk of positivity bias, where users tend to share mostly positive aspects of their experiences [32]. Another important source for mental health research is clinical data, which typically contains sensitive information and is not publicly accessible.
Most research on mental health is conducted using data in English, which may overlook important nuances present in other languages. Previous studies suggest that features predictive of mental health issues in English may not translate effectively to other languages or cultures [8, 30]. For example, the use of the first-person pronoun "I" has been identified as a strong predictor of depression [31], and this feature is frequently analyzed in English-based research. However, the relationship between pronoun usage and depression severity can vary significantly across different demographic groups and languages [30]. While the pronoun "I" is often used as an indicator of depression in English, its usage requires special consideration in other languages due to linguistic differences. For instance, previous research has shown that this association has not been observed among Romanian speakers [36]. Romanian is a pro-drop language, which allows for ambiguity in subject pronoun usage and may contribute to a lower frequency of the personal pronoun ”I” in everyday communication.
Despite the availability of some data in the Romanian language related to emotions, such as the RED dataset [10], REDv2 [9], and the BRIGHTER multilingual dataset [27], which includes Romanian data, there is a scarcity of mental health-related resources. Very few clinical data sources [36] and posts from medical forums [7, 12] exist, and those that do are not publicly accessible.
This study aims to address this gap by creating the first publicly available dataset in Romanian, specifically designed to analyze emotional expressions in the context of depression and anxiety. By analyzing the linguistic characteristics and emotional patterns in Romanian, our research will contribute to a deeper understanding of mental health expressions in languages other than English. Through this initiative, we seek to promote more inclusive research in mental health, paving the way for targeted interventions and support across diverse linguistic communities.
2 Related Work
Previous research has gathered open-ended responses from participants to explore how expressions of mental health influence language. [39] created open-ended personality questions to explore how natural language processing can aid in predicting neuroticism from people’s responses. They gathered over 30,000 sentences written by 425 adults in Korean and used KoBERT to see which types of questions best revealed neuroticism and related traits like depression and dependency. [37] created the multi-modal Real World Worry Waves Dataset (RW3D), in English, which tracks the psychological effects of the COVID-19 pandemic in the UK over three consecutive years, 2020, 2021 and 2022, through repeated surveys containing also free-text responses. [11] used open-ended interview responses from more than 58,000 job candidates along with self-assessed personality (HEXACO) and employed “InterviewBERT” to deduce personality traits from textual replies. This study was also done in English. [33] used open-ended questionnaires to study young and middle aged adults’ responses regarding their own mental health, by employing NLP computational methods. They analyzed their 1106 responses and found various differences in semantics on the two study groups.
In terms of combining emotion detection and LIWC term analysis in mental health texts, several studies employ such analysis in English: [29], [40], [2]. [4] combined LIWC features [6], sentiment analysis and NLP to explore the relationship between language and mental health issues such as depression, anxiety, and stress during the COVID-19 pandemic. They analyzed written narratives from 883 participants, alongside their DASS-21 scores, and found an association between increased symptom severity and greater use of first-person pronouns and negative emotion words. [21] collected responses in English from 411 participants who answered word-response questions related to their mental health and completed the GAD-7 and PHQ-9 rating scales. Their analyses show that individuals’ language reflects their mental health status, and combining verbal responses with quantitative measures can enhance the comprehension and assessment of the risk of depressive and anxiety symptoms in both clinical practice and research contexts.
While several datasets have been created using text gathered from open-ended questionnaires, combined with screening tools to assess the risk of mental health issues, most of these resources are focused on the English language. Currently, there are no open-source psychological corpora available in Romanian, and this was the main driving force behind our study, along with our interest in applying emotional analysis to such a corpus.
3 Data Collection
As opposed to social media collected data, which is collected without participants’ consent, we collected data from participants who provided their informed consent for data collection [5]. The data collection methodology consists of a completely anonymous survey, organized in three parts. The first part of the survey consists of six open-ended questions, designed to gather extensive written responses. Out of these six questions, three touch on positive topics and three on negative topics. The survey was written in Romanian for Romanian speakers only. We further present the six open-ended questions, in both English and Romanian:
  1. 1.
    To what extent do you think social media affects your emotional state, especially when you see negative news? “În ce măsură crezi că rețelele sociale îți afectează starea emoțională, mai ales când vezi știri negative?” (negative)
  2. 2.
    What is an important lesson about resilience and adaptation that you have learned recently? “Care este o lecție importantă despre reziliență și adaptare pe care ai învățat-o recent?” (positive)
  3. 3.
    How much do you think politics affects people’s mental health? Can you describe a time when you felt stressed because of political decisions? “Cât de mult crezi că politica afectează sănătatea mentală a oamenilor? Poți să detaliezi un moment în care ai simțit stres din cauza deciziilor politice?” (negative)
  4. 4.
    What role do you think community or social support plays in maintaining mental health during difficult times? “Ce rol crezi că are comunitatea sau sprijinul social în menținerea sănătății mentale în vremuri dificile?” (positive)
  5. 5.
    How were you affected by discussions about war, the economic crisis, elections, or other tense topics? “Cum ai fost afectat de discuțiile despre război, criza economică, alegeri sau alte subiecte tensionate?” (negative)
  6. 6.
    How do you recharge your batteries after a stressful period? “Cum îți reîncarci bateriile după o perioadă stresantă?” (positive)
Usually, psychological corpora for NLP are gathered from social media through searching for different keywords in online posts that could indicate depression and anxiety. However, after consulting with medical professionals in the psychiatric field, we have concluded that this method of collecting psychological data may not be entirely accurate. Thus, we added two more parts to our survey: one assessing respondents’ risk of depression using the PHQ-9 [23], and another assessing their risk of anxiety using the GAD-7 [35]. Both GAD-7 and PHQ-9 contain questions (7 for GAD-7 and 9 for PHQ-9) that ask the participants how often they have experienced specific problems, and the response options correspond to a numerical score: Not at all = 0, Several days = 1, More than half the days = 2, Nearly every day = 3. The severity cut-offs for depression based on PHQ-9 are: 0-4: Minimal, 5-9: Mild, 10-14: Moderate, 15-19: Moderately severe, 20-27: Severe, while the severity cut-offs for anxiety based on GAD-7 are: 0-4: Minimal, 5-9: Mild, 10-14: Moderate, 15-19: Moderately severe, 20-27: Severe.
The resulting questionnaire is completely anonymous, and no personal data is requested or stored (such as name, surname, email address, or IP). We did not collect demographic data for the same reason: to build trust with respondents and keep the time required to complete the form short.
The resulting data consists of 205 responses, collected over an eight-month period from March 2025 to October 2025. This time frame includes two major stressful events for Romanians - presidential elections and the tax increase burden generated by the country’s new leadership.
4 Methodology
Our research consists of gathering insights from PsihoRo descriptive statistics in correlation to depression and anxiety respondents’ tendencies, for three main population groups which were computed based on the severity cut-offs. According to the literature in psychology [22], a cut-off score of 10 or higher on the PHQ-9 questionnaire is used to identify the presence of potential major depressive disorder, prompting further assessment. For the GAD-7 questionnaire, also a cut-off score of 10 has been proposed as a threshold for identifying potential generalized anxiety disorder, warranting further evaluation. [34].
Based on these rules, three population groups were derived from our dataset:
  • Depression Risk Group: 37 respondents scored higher than 10 on PHQ-9 (18%);
  • Anxiety Risk Group: 36 respondents scored higher than 10 on GAD-7 (17.56%);
  • No Risk Group: 28 respondents scored 0 for at least one out of the two questionnaires (13.66%).
Linguistic Inquiry and Word Count (LIWC) [6] is a lexicon and a text analysis software that provides psychologically meaningful categories and is widely used in computational mental health research [31, 24, 17, 36]. The official LIWC software is available for analyzing textual data in the English language; however, validated translations of the software are also available in several languages.
In our experiments, we use the Romanian version of LIWC (Ro-LIWC2015) [12]. The 2015 version of LIWC includes 98 categories that relate to linguistic dimensions and grammar, as well as psychological processes, including affective, cognitive, social, perceptual, and biological processes.
Previous research has indicated that individuals with mental health issues tend to exhibit specific language patterns. These patterns include a higher frequency of using the personal pronoun "I" [31], increased use of words with negative connotations [14], and a tendency to use words with absolute meanings [1]. These findings highlight the importance of language analysis in understanding mental health and motivate us to use LIWC to examine the language patterns of the participants. After checking for differences in LIWC categories in relation to the PHQ-9 and GAD-7 scores, we wanted to test their discriminatory power in classification. To this end, we employed a LightGBM classifier [20] and trained it on the results with LIWC to predict a binary label for depression and anxiety, based on the cutoff score of 10. The model achieved an accuracy of 0.85 in classifying depression and 0.83 in classifying anxiety. To determine which LIWC features have the greatest influence on predictions, we used SHapley Additive exPlanations (SHAP) [26].
To complement these analyses, we have created an emotion detection model by fine-tuning the Romanian BERT [13] on the REDv2 dataset [9], for multi-label text classification, using the Huggingface transformers pipeline [^1] https://huggingface.co/docs/transformers. We used this model to predict the emotions in PsihoRo, thus gathering emotional information for the text written by each population group mentioned above. The resulting model had an F1 score of 66.85% computed on the REDv2 test set.
Another interesting analysis was topic modeling. To identify the main themes in the PsihoRo corpus we used the stm R package. First, the texts were tokenized, lowercased, and cleaned by removing punctuation, numbers, symbols and stopwords. Next, we created a Document-Feature Matrix (DFM) to represent term frequencies across texts. Also, to enhance model performance and reduce noise, we excluded rare terms which appeared in fewer than 3 documents and the very common terms, which appeared in more than 90% of the texts. Further, the cleaned DFM was transformed into the format necessary for the Structural Topic Model (STM) using which we trained the model for K = 5 topics, with the Spectral initialization method.
5 Experiments and Results
5.1 PsihoRo Descriptive Statistics
In Table 1 we show mean, median (50% quantile), min, max, 25% and 75% cuantiles, and the standard deviation for both anxiety and depression scores. It can be seen that the maximum value for GAD-7 (anxiety) is 21, which is the maximum which can be achieved, and the maximum score for PHQ-9 (depression) is 23, out of the maximum possible of 27.
Table 1: Descriptive statistics for PsihoRo
PHQ-9 GAD-7
count 205 205
mean 6 5.78
standard deviation 4.83 4.95
min 0 0
25% 3 2
50% (median) 5 5
75% 8 7
max 23 21
We analyzed the length of responses to determine whether individuals at risk of depression or anxiety tend to write less. Our findings indicate that this is not the case for Romanian. The computed T-statistic was 1.55 when comparing the depression risk group with the group not at risk, and 2.26 when comparing the anxiety risk group with the group not at risk. These results suggest that there is no significant difference in response length between these populations.
In Figure 1 we show the distribution plots of both PHQ-9 and GAD-7 scores. The data indicate that the distributions are skewed to the left, with the majority of observations concentrated at lower values and a longer tail extending toward the higher end of the scale, meaning that most respondents have minimal and mild risk of depression and anxiety.
Figure 1: Distributions for GAD-7 and PHQ-9
The Pearson correlation coefficient (r) between PHQ-9 and GAD-7 total score is 0.761 (p<0.001), which means that anxiety and depression are strongly correlated in our data. This result shows as well in Figure 2 and is in concordance with the results of several other works: [19], [16], [18].
Figure 2: Correlation between Anxiety and Depression Scores.
5.2 LIWC
We employ LIWC to analyze language patterns in participants’ open-ended responses and correlate them with scores from the PHQ-9 and GAD-7. In Table 2, we present the Pearson correlation coefficients between LIWC categories and the symptom scores. Positive values indicate that higher symptom scores are associated with greater use of that linguistic category, while negative values indicate the opposite. Asterisks indicate significance (*p < .05, **p < .01).
Surprisingly, the use of negations is negatively correlated with both PHQ-9 and GAD-7 scores, suggesting that participants with higher depression and anxiety scores tend to use fewer negations in their responses. Individuals who score higher on these assessments tend to use more words in the Tentative category, which indicates uncertainty, supported by the correlation coefficients of .204** for PHQ-9 and .213** for GAD-7. Comparatives also show a positive correlation with the depression score (.167*), indicating comparative thinking, which is often linked to depressive rumination. However, another marker for rumination, the high frequency of Past focus verbs, is not observed in our data. There is a positive correlation between the use of Body-related words and PHQ-9 (200**) and GAD-7 (180**) scores, indicating that participants with higher scores frequently use more language related to bodily sensations. This may reflect increased somatic awareness, a common trait in depression and anxiety [28]. Participants with higher PHQ-9 and GAD-7 scores discuss Achievement less frequently (-.163* and- .192**, respectively), which could indicate low motivation. In addition, the negative correlation with Leisure-related words (-.145* and -.169*, respectively) may stem from anhedonia, a common symptom of depression ([15]). Words in the Death category show a significant positive correlation with depression (.138*), but not with anxiety (0.029), indicating a greater preoccupation with mortality in participants with higher depression scores. In contrast, the Family category is positively correlated with anxiety scores (r = 0.145*), suggesting that individuals with higher anxiety are more focused on family-related issues.
Table 2: Correlations between PHQ-9, GAD-7, and LIWC dimensions
LIWC Dimension Examples PHQ-9 GAD-7
I I, me, mine 0.018 -0.046
We we, us, our 0.013 -0.039
You you, your 0.006 0.084
Verb eat, come, carry -0.087 -0.091
Past focus ago, did, talked -0.064 -0.075
Present focus today, is, now -0.019 -0.034
Future focus may, will, soon -0.027 -0.055
Adjectives free, happy, long 0.118 0.12
Negations no, not, never -.156* -.167*
Positive emotion love, nice, sweet -0.122 -0.057
Negative emotion hurt, ugly, nasty 0.028 0.028
Anxiety worried, fearful -0.042 0.01
Anger hate, kill, annoyed 0.107 0.105
Sadness crying, grief, sad 0.074 0.045
Family daughter, dad, aunt 0.079 .145*
Friends buddy, neighbor -0.026 0.043
Female references girl, her, mom 0.061 0.074
Male references boy, his, dad .156* .177*
Comparatives greater, best, after .167* 0.081
Causation because, effect -.186** -.180**
Tentative maybe, perhaps .204** .213**
Insight think, know -0.075 -0.101
Assent agree, OK, yes 0.009 0.065
Perceptual processes look, heard, feeling 0.009 0.033
Body cheek, hands, spit .200** .180**
Health clinic, flu, pill -0.051 0.035
Achievement win, success, better -.163* -.192**
Leisure cook, chat, movie -.145* -.169*
Death bury, coffin, kill .138* 0.029
Note. * $p<.05$, ** $p<.01$.
Figure 3: Differences in selected LIWC linguistic categories between participants with low vs. high PHQ-9 (left) and GAD-7 (right) scores. * denotes statistically significant differences ($p<0.05$) measured using the Mann-Whitney U test. Positive values indicate greater relative use among participants with higher symptom scores.
To better understand the differences between higher and lower symptom scores for depression and anxiety, we conducted an additional analysis of the LIWC categories. According to the cutoff values for both the PHQ-9 and GAD-7 assessment tools, we define a low PHQ score as anything less than 10, while a score greater than 10 is classified as a high PHQ-9 score. In Figure 3, the goal was to identify linguistic markers that differentiate individuals with higher vs. lower symptom scores. Statistically significant differences are indicated with an asterisk (*), as determined by the Mann-Whitney U test. Positive values indicate greater relative use among participants with higher symptom scores, while negative values indicate greater use among low-symptom participants. The analysis produced results consistent with the correlation analysis, particularly regarding Achievement and tentativeness in relation to depression and anxiety. However, for depression, we observed a decreased use of Positive emotion words and an increased usage of Sadness words, which aligns with existing literature in psychology [36]. In addition, the language of participants with high GAD-7 scores showed a higher frequency of words in the Interrogatives category, indicating a tendency to question or seek clarity.
Figure 4: Feature importance for LIWC categories analyzed using SHAP.
To determine which LIWC features have the greatest influence on predictions, we used SHapley Additive exPlanations (SHAP) [26], as explained in Section 4, with the results presented in Figure 4. The importance of the top 20 features aligns with our correlation analysis and highlights the differences between low and high scores for both depression and anxiety.
The LIWC analyses presented in this section demonstrate that lexical indicators can effectively capture underlying psychological patterns related to depression and anxiety. However, the most important feature used in mental health research, the pronoun ”I”, appears to be less useful for the Romanian language, likely because it is a pro-drop language. This finding is consistent with the work of [36], which also indicated that this marker does not generalize well in a sample of patients with major depression.
5.3 Emotion Analysis
Using the emotion detection model described in Section 4, we predicted the emotions in all the texts of PsihoRo, and then we aggregated the results by using radar charts to visualize the overall emotional patterns. In Figure 5 we show the emotion predictions across all PsihoRo texts, and also for the pessimistic questions (1, 3, 5) and the optimistic ones (2, 4, 6).
Figure 5: Emotion predictions across: all PsihoRo texts (left), pessimistic questions (middle), and optimistic questions (right).
Figure 6: Emotion predictions across the three studied population groups.
Surprise is an emotion that is completely absent from the PsihoRo data. The following emotions are more common, in order of appearance: anger, joy, fear, sadness, trust, and the most common, neutral. The emotional pattern for the more pessimistic questions is displayed in the center of Figure 5, while the emotional pattern for the more optimistic questions is displayed on the right. Emotions such as sadness and anger are on the rise on the Pessimistic Questions, while joy and trust are more prevalent on the Optimistic Questions.
The emotions present in the texts written by the three primary demographic groups mentioned at the beginning of Section 4 are presented in Figure 6. While the neutral emotion is strongly represented across all groups, the population at risk for depression shows a high prevalence of sadness. In addition, the group at risk for anxiety exhibits emotions of sadness, fear and trust.
5.4 Topic Modeling
We applied the topic modeling analysis described in Section 4 for each of the three population groups studied in this work. The main topics are derived from the six open questions asked in the form. However, some differences appear among the three groups of respondents, as discussed in the following subsections.
5.4.1 No Risk Group
Topic 1 highlights how individuals perceive the current political and social period in a critical and sometimes negative way. The words “negative”, “decisions”, “politics”, “measure”, “news” and “period” indicate a concern for the effects of government decisions and how they are reflected in the media. The presence of the term “sport” suggests a possible comparison between different areas of public life.
Topic 2 reflects the complex interaction between information conveyed through news, political life, and how they influence the general state of the population. Terms such as “politics”, “news”, “state”, “people” and “politics” indicate an increased attention to how media discourse shapes collective perceptions and emotions.
Topic 3 addresses how individuals interpret and process information related to elections and political decisions. The terms “measure”, “depends”, “people”, “time”, “elections” and “example” suggest a personal analysis of the electoral context and an attempt to understand the relationship between information, time and decision.
Topic 4 highlights the importance of authentic human interaction and the balance between the inner and social dimensions of the individual. The words “listen”, “outer”, “inner”, “community” and “matter” suggest an openness to empathetic communication and highlight the value of social belonging.
Topic 5 focuses on the emotional dimension of everyday experiences, indicating how individuals manage their affective states and reactions in common situations. Words such as “affect”, “try”, “day”, “go” and “enough” express a constant effort to adapt and maintain psychological balance in the face of daily pressures by self-regulation and personal reflection.
5.4.2 Depression Risk Group
Topic 1 captures the perception of fatigue, daily routine and the constant effort that individuals put into their personal and professional lives. Associated terms such as “sleep”, “hard”, “walking” and “people” indicate a general state of exhaustion and the need for balance between personal and social life.
Topic 2 highlights the importance of psychological resilience and the ability to adapt to stressful or unpredictable situations. Words such as “resilience”, “situation”, “time”, “community” and “important” suggest a solution-oriented approach and maintaining a state of balance in contexts of uncertainty.
Topic 3 focuses on how individuals relate to political processes and the electoral climate. The terms “politics”, “elections”, “people”, “moment” and “listening” indicate both interest in political events and their influence on social attitudes and conversations.
Topic 4 captures the potentially harmful impact of social networks on interpersonal relationships and community cohesion. The presence of the words “negative”, “social”, “networks”, “community” and “socialization” underlines the perception of a degradation of authentic communication and well-being caused by the digital environment.
Topic 5 reflects the concern for affective states and how social and economic factors contribute to the intensification of collective anxiety. The terms “feel”, “affect”, “cause”, “anxiety”, “country” and “family” suggest an emotional perspective on everyday experiences and the general social context.
5.4.3 Anxiety Risk Group
Topic 1 captures how major events, such as war, economic crises or socio-political changes, affect the perception of daily life. Words such as “affect”, “crisis”, “war”, “elections”, “life” and “time” indicate the concern for external instability and its influence on personal balance.
Topic 2 highlights the close connection between political factors and individual emotional state. Terms such as “politics”, “stress”, “affect”, “feel”, “cause” and “family” suggest a direct relationship between socio-political events and the level of anxiety or psychological tension felt.
Topic 3 addresses the role of community and collective resilience in the face of the constant flow of negative information from the media. The words “resilience”, “news”, “negative”, “community”, “important” and “trying” indicate an attempt to find the balance between information and protecting one’s mental state.
Topic 4 explores the effects of social media on psychological state, with a focus on anxiety and social pressure. Terms such as “social”, “anxiety”, “state”, “hard”, “networks” and “present” suggest a connection between the intensive use of social networks and the degradation of emotional state. The topic reflects how virtual interactions can accentuate psychological discomfort.
Topic 5 emphasizes the connection between people, community and the sense of belonging. The words “people”, “community”, “listen”, “dear” and “happen” highlight the importance of close relationships and solidarity in maintaining emotional balance.
6 Conclusion and Future Work
In this paper, we presented PsihoRo, the first open-source dataset for mental health in Romanian. The dataset was collected from 205 participants who filled in a comprehensive form that included both open-ended questions for text collection and validated questionnaires, specifically the PHQ-9 and GAD-7, to evaluate their risk of depression and anxiety.
We analyzed the collected data using the psychologically meaningful categories from LIWC, emotion analysis and topic modeling. We observed differences in LIWC categories related to body-related words, which may indicate somatic awareness, and tentative words, which indicate uncertainty, that correlated positively with both depression and anxiety scores. A higher preoccupation with mortality was also observed in relation to depression, while anxiety scores were positively correlated with a preoccupation with family. While LIWC analyses showed that lexical indicators can effectively capture underlying psychological patterns related to depression and anxiety, we did not find any association between the linguistic marker “I” and depression or anxiety in the Romanian language, consistent with previous research [36]. The emotion analysis revealed a prevalence of sadness in the population group at risk for depression, as well as fear in the population group at risk for anxiety, which aligns with previous research [25, 3, 38]. Topic modeling showed that although some perceptions and emotions are captured in textual responses, a multi-perspective approach that goes beyond self-expressed text is necessary to fully comprehend these patterns. By offering high-quality data with precise labels for anxiety and depression risk, PsihoRo seeks to supplement this intricate research.
7 Ethical Considerations and Limitations
The data collection received approval from the Ethics Committee of the University of Bucharest. All respondents provided their informed consent before participating in the survey, which was completely anonymous. No personal information, such as names, surnames, email addresses, or IP addresses, was requested or stored.
We believe that increasing the number of respondents through crowdsourcing or similar platforms could significantly enhance PsihoRo. We aim to implement this in future versions of this resource.
Acknowledgments
The authors would like to thank psychologist Virginia Neacșu for her valuable feedback and suggestions throughout the development of this work.
This work was partially supported by Druid AI, University of Bucharest and the Ministry of Education and Research, CNCS-UEFISCDI, project SIROLA, number PN-IV-P1- PCE-2023-1701, within PNCDI IV, and by the project "Romanian Hub for Artificial Intelligence - HRIA”, Smart Growth, Digitization, and Financial Instruments Program, 2021-2027, MySMIS no. 334906.